# LLMs for Knowledge Graph Construction and Reasoning: Recent Capabilities and Future Opportunities

### v2

@@ 2.3 Discussion: Why LLMs do not present satisfactory performance on some tasks?
Our experiments underscore GPT-4’s ability to extract knowledge across diverse domains, albeit not surpassing the performance of fine-tuned models. This observation also aligns with findings from previous research  (Wei et al., 2023b; Gao et al., 2023). Our experiment, conducted in March-April 2023, uses an interactive interface rather than an API to evaluate the GPT models on a randomly selected subset of datasets.
Notably, in assessing large models across eight datasets, we identify that the outcomes may be subject to various factors. Dataset Quality: Using the KG construction task as an illustration, dataset noise could lead to ambiguities. Complex contexts and potential label inaccuracies may also negatively impact model evaluation. Instruction Quality: Model performance is notably influenced by the semantic depth of instructions. Finding optimal instructions through prompt engineering  [^3] https://www.kdnuggets.com/publications/sheets/ChatGPT_Cheatsheet_Costa.pdf can enhance performance. An In-context Learning (Dong et al., 2023) approach with relevant samples can further improve outcomes. Evaluation Methods: Current methods may not be entirely apt for assessing the capabilities of large models like ChatGPT and GPT-4. Dataset labels may not capture all correct responses, and answers involving synonymous terms might not be accurately recognized.
@@ 2.4 Discussion: Do LLMs have memorized knowledge or truly have the generalization ability?
To explore this, we design the Virtual Knowledge Extraction task, targeting LLMs’ ability to generalize and extract unfamiliar knowledge. Given the insufficiency of existing datasets, we present VINE, a new dataset designed specifically for Virtual Knowledge Extraction.
@@ 2.4.2 Preliminary Results
In Figure 5, we provide large models with a triple composed of virtual relation types and virtual head and tail entities—[Schoolnogo, decidiaster, Reptance] and [Intranguish, decidiaster, Nugculous]—along with the respective demonstrations. The results demonstrate that GPT-4 effectively completed the extraction of the virtual triple. Consequently, we tentatively conclude that GPT-4 exhibits a relatively strong generalization ability and can rapidly acquire the capability to extract new knowledge through instructions, rather than relying solely on the memory of relevant knowledge. Related work (Wei et al., 2023a) has also confirmed that large models possess an exceptionally strong generalization ability concerning instructions.
@@ 3 Future Opportunities: Automatic KG Construction and Reasoning
Consequently, there is a growing interest in the realm of interactive natural language processing (iNLP) (Wang et al., 2023b). In parallel, research efforts concerning intelligent agents continue to proliferate Wang et al. (2023a); Xi et al. (2023); Zhao et al. (2023a). A notable example of this advancement is AutoGPT [^4] https://github.com/Significant-Gravitas/Auto-GPT, which can independently generate prompts and carry out tasks such as event analysis, programming, and mathematical operations. Concurrently, Li et al. (2023) delves into the potential for autonomous cooperation between communicative agents and introduces a novel cooperative agent framework called role-playing.
@@
For a simple demonstration of the concept, we utilize the role-playing method in CAMEL (Li et al., 2023). As depicted in Figure 6, we designate the KG assistant agent as a Consultant and the KG user agent as a KG domain expert. Upon receipt of the prompt and assigned roles, the task-specifier agent provides an elaborate description to clarify the concept. Following this, the KG assistant and KG user collaborate in a multi-party setting to complete the specified task until the KG user confirms its completion. Concurrently, a web searcher role is introduced to aid the KG assistant in internet knowledge retrieval. When the KG assistant receives a dialogue from the KG user, it initially consults the web searcher on whether to browse information online based on the content. Guided by the web searcher’s response, the KG assistant then continues to address the KG user’s command. The experimental example indicates that the knowledge graph related to the film Spider-Man: Across the Spider-Verse released in 2023 is more effectively and comprehensively constructed using the multi-agent and internet-augmented approach.
Remark.
By combining the efforts of artificial intelligence and human expertise, AutoKG could speed up the creation of specialized KGs, fostering a collaborative environment with language models. This system leverages domain and internet knowledge to produce high-quality KGs, augmenting the factual accuracy of LLMs in domain-specific tasks, thereby increasing their practical utility. AutoKG not only simplifies the construction process but also improves LLMs’ transparency, facilitating a deeper understanding of their internal workings. As a cooperative human-machine platform, it bolsters the understanding and guidance of LLMs’ decision-making, increasing their efficiency in complex tasks. However, it is noteworthy that despite the assistance of AutoKG, the current results of the constructed knowledge graph still necessitate manual evaluation and validation.
@@ 4 Conclusion and Future Work
Limitations
While our research has yielded some results, it also possesses certain limitations. As previously stated, the inability to access the GPT-4 API has necessitated our reliance on an interactive interface for conducting experiments, undeniably inflating workload and time costs. We look forward to future research opportunities that will allow us to further explore these areas.
LLMs.
We confine our experiments to models within the GPT series, leaving the performance of other large models like LaMDA (Thoppilan et al., 2022) unexamined. Future work could extend these experiments to more LLMs. Additionally, we do not have access to the GPT-4 API; thus, we complete our experiments via an interactive interface, which is both time-consuming and labor-intensive.
Tasks.
Not all KG construction and reasoning tasks are considered in our study. We focus on a handful of representative tasks, which might limit the applicability of our findings in specific contexts. Also, due to the unavailability of GPT-4’s multimodal capabilities to the public, we are unable to delve into its performance and contribution to multimodal processing. We look forward to future research opportunities that would allow us to explore these areas further.
Ethical and Factual Considerations
Large language models used in our experiments may have inherent biases and issues related to factual accuracy. Thus, the experimental results should be interpreted with a critical mindset.
@@ A.1 Large Language Models
LLMs are pre-trained on substantial amounts of textual data and have become a significant component of contemporary NLP research. Recent advancements in NLP have led to the development of highly capable LLMs, such as GPT-3 (Brown et al., 2020), ChatGPT, and GPT-4, which exhibit exceptional performance across a diverse array of NLP tasks, including machine translation, text summarization, and question answering. Concurrently, several previous studies have indicated that LLMs can achieve remarkable results in relevant downstream tasks with minimal or even no demonstration in the prompt (Wei et al., 2022; Bang et al., 2023; Nori et al., 2023; Qiao et al., 2023; Wei et al., 2023b). This provides further evidence of the robustness and generality of LLMs.
@@ A.2 ChatGPT & GPT-4
ChatGPT, an advanced LLM developed by OpenAI, is primarily designed for engaging in human-like conversations. During the fine-tuning process, ChatGPT utilizes RLHF (Christiano et al., 2017), thereby enhancing its alignment with human preferences and values.
@@
With the popularity of LLMs, an increasing number of researchers are exploring the specific emergent capabilities and advantages they possess. Bang et al. (2023) performs the in-depth analysis of ChatGPT on the multitask, multilingual and multimodal aspects. The findings indicate that ChatGPT excels at zero-shot learning across various tasks, even outperforming fine-tuned models in certain cases. However, it faces challenges when generalized to low-resource languages. Furthermore, in terms of multi-modality, compared to more advanced vision-language models, the capabilities of ChatGPT remain fundamental. Moreover, ChatGPT has garnered considerable attention in other various domains, including information extraction (Gao et al., 2023; Wei et al., 2023b), reasoning (Qin et al., 2023), text summarization (Jeblick et al., 2022), question answering (Tan et al., 2023) and machine translation (Jiao et al., 2023), showcasing its versatility and applicability in the broader field of natural language processing.
While there is a growing body of research on ChatGPT, investigations into GPT-4 continue to be relatively limited. Nori et al. (2023) conducts an extensive assessment of GPT-4 on medical competency examinations and benchmark datasets and shows that GPT-4, without any specialized prompt crafting, surpasses the passing score by over 20 points. Kasai et al. (2023) also studies GPT-4’s performance on the Japanese national medical licensing examinations. Furthermore, there are some studies on GPT-4 that focus on cognitive psychology (Sifatkaur et al., 2023), academic exams (Nunes et al., 2023), and translation of radiology reports (Lyu et al., 2023).
@@ Appendix B Datasets
Entity, Relation and Event Extraction.
DuIE2.0 (Li et al., 2019) is a substantial Chinese relationship extraction dataset with more than 210,000 sentences and 48 predefined relationship categories. SciERC (Luan et al., 2018) is a collection of scientific abstracts annotated with seven relations. Re-TACRED (Stoica et al., 2021), an upgraded version of the TACRED dataset, includes over 91,000 sentences across 40 relations. MAVEN (Wang et al., 2020) a general-domain event extraction benchmark containing 4,480 documents and 168 event types.
Link Prediction
FB15K-237 (Toutanova et al., 2015) is widely used as a benchmark for assessing the performance of knowledge graph embedding model on link prediction, encompassing 237 relations and 14,541 entities. ATOMIC 2020 (Hwang et al., 2021a) serves as a comprehensive commonsense repository with 1.33 million inferential knowledge tuples about entities and events.
Question Answering
FreebaseQA (Jiang et al., 2019) is an open-domain QA dataset built on the Freebase knowledge graph, comprising various sourced question-answer pairs. MetaQA (Zhang et al., 2018), expanded from WikiMovies (Miller et al., 2016), provides a substantial collection of single-hop and multi-hop question-answer pairs, surpassing 400,000 in total.
@@ Appendix C Data Collection of VINE
Using GPT-4 data up to September 2021 as a basis, we select a portion of participants’ responses from two competitions organized by the New York Times as part of our data sources. These competitions include the "February Vocabulary Challenge: Invent a Word" [^5] https://www.nytimes.com/2022/01/31/learning/february-vocabulary-challenge-invent-a-word.html held in January 2022 and the "Student Vocabulary Challenge: Invent a Word" [^6] https://www.nytimes.com/2023/02/01/learning/student-vocabulary-challenge-invent-a-word.html conducted in February 2023. Both competitions aim to promote the creation of distinctive and memorable new words that address gaps in the English language.
@@ Appendix D Prompts for Evaluation
Event Detection
The list of event types: [……]
@@
Link Prediction
predict the tail entity [MASK] from the given (40th Academy Awards, time event locations, [MASK]) by completing the sentence "what is the locations of 40th Academy Awards? The answer is".
predict the tail entity [MASK] from the given (1992 NCAA Men’s Division I Basketball Tournament, time event locations, [MASK]) by completing the sentence "what is the locations of 1992 NCAA Men’s Division I Basketball Tournament? The answer is ".The answer is Albuquerque, so the [MASK] is Albuquerque.
predict the tail entity [MASK] from the given (40th Academy Awards, time event locations, [MASK]) by completing the sentence "what is the locations of 40th Academy Awards? The answer is".
Question Answering
@@
Answer: Android | City Limits

### v3

@@ 2.3 Discussion: Why LLMs do not present satisfactory performance on some tasks?
Our experiments underscore GPT-4’s ability to extract knowledge across diverse domains, albeit not surpassing the performance of fine-tuned models. This observation also aligns with findings from previous research  [25, 53]. Our experiment, conducted in March-April 2023, uses an interactive interface rather than an API to evaluate the GPT models on a randomly selected subset of datasets.
Notably, in assessing large models across eight datasets, we identify that the outcomes may be subject to various factors. Dataset Quality: Using the KG construction task as an illustration, dataset noise could lead to ambiguities. Complex contexts and potential label inaccuracies may also negatively impact model evaluation. Instruction Quality: Model performance is notably influenced by the semantic depth of instructions. Finding optimal instructions through prompt engineering  [^3] https://www.kdnuggets.com/publications/sheets/ChatGPT_Cheatsheet_Costa.pdf can enhance performance. An In-context Learning [54] approach with relevant samples can further improve outcomes. Evaluation Methods: Current methods may not be entirely apt for assessing the capabilities of large models like ChatGPT and GPT-4. Dataset labels may not capture all correct responses, and answers involving synonymous terms might not be accurately recognized.
@@ 2.4 Discussion: Do LLMs have memorized knowledge or truly have the generalization ability?
To explore this, we design the Virtual Knowledge Extraction task, targeting LLMs’ ability to generalize and extract unfamiliar knowledge. Unlike conventional benchmarks, the task focuses on evaluating how models perform when confronted with information they have not previously encountered, rather than relying solely on knowledge accumulated during pre-training. Existing datasets largely comprise entities familiar to LLMs, potentially sourced from their pre-training corpora, thereby possibly including relationships already encoded within these corpora during extraction tasks. Addressing these dataset constraints, we introduce VINE, a novel dataset specifically crafted for Virtual Knowledge Extraction.
@@ 2.4.2 Preliminary Results
In Figure 4, we provide large models with a triple composed of virtual relation types and virtual head and tail entities—[Schoolnogo, decidiaster, Reptance] and [Intranguish, decidiaster, Nugculous]—along with the respective demonstrations. The results demonstrate that GPT-4 effectively completed the extraction of the virtual triple. Consequently, we tentatively conclude that GPT-4 exhibits a relatively strong generalization ability and can rapidly acquire the capability to extract new knowledge through instructions, rather than relying solely on the memory of relevant knowledge. Related work [55] has also confirmed that large models possess an exceptionally strong generalization ability concerning instructions.
@@ 3 Future Opportunities: Automatic KG Construction and Reasoning
Consequently, there is a growing interest in the realm of interactive natural language processing (iNLP) [37]. In parallel, research efforts concerning intelligent agents continue to proliferate [56, 57, 58]. A notable example of this advancement is AutoGPT [^4] https://github.com/Significant-Gravitas/Auto-GPT, which can independently generate prompts and carry out tasks such as event analysis, programming, and mathematical operations. Concurrently, Li et al. [59] delves into the potential for autonomous cooperation between communicative agents and introduces a novel cooperative agent framework called role-playing.
@@
For a simple demonstration of the concept, we utilize the role-playing method in CAMEL [59]. As depicted in Figure 5, we designate the KG assistant agent as a Consultant and the KG user agent as a KG domain expert. Upon receipt of the prompt and assigned roles, the task-specifier agent provides an elaborate description to clarify the concept. Following this, the KG assistant and KG user collaborate in a multi-party setting to complete the specified task until the KG user confirms its completion. Concurrently, a web searcher role is introduced to aid the KG assistant in internet knowledge retrieval. When the KG assistant receives a dialogue from the KG user, it initially consults the web searcher on whether to browse information online based on the content. Guided by the web searcher’s response, the KG assistant then continues to address the KG user’s command. The experimental example indicates that the knowledge graph related to the film Spider-Man: Across the Spider-Verse released in 2023 is more effectively and comprehensively constructed using the multi-agent and internet-augmented approach.
Remark. By combining the efforts of artificial intelligence and human expertise, AutoKG could speed up the creation of specialized KGs, fostering a collaborative environment with language models. This system leverages domain and internet knowledge to produce high-quality KGs, augmenting the factual accuracy of LLMs in domain-specific tasks, thereby increasing their practical utility. AutoKG not only simplifies the construction process but also improves LLMs’ transparency, facilitating a deeper understanding of their internal workings. As a cooperative human-machine platform, it bolsters the understanding and guidance of LLMs’ decision-making, increasing their efficiency in complex tasks. However, it is noteworthy that despite the assistance of AutoKG, the current results of the constructed knowledge graph still necessitate manual evaluation and validation.
@@
We also notice some existing works related to KG con
@@ 4 Conclusion and Future Work
Declarations
  • Funding. This work was supported by the National Natural Science Foundation of China (No. 62206246, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China (No. LGG22F030011), Yongjiang Talent Introduction Programme (2021A-156-G), Tencent AI Lab Rhino-Bird Focused Research Program (RBFR2024003), Information Technology Center and State Key Lab of CAD&CG, Zhejiang University, and NUS-NCS Joint Laboratory (A-0008542-00-00).
  • Ethics approval and consent to participate. This work did not involve any human participants, their data, or biological materials, and therefore did not require ethical approval.
  • Data and Materials availability. Our data and materials are accessible in the repository here  [^5] The code and datasets are in https://github.com/zjunlp/AutoKG.
@@ A.1 Large Language Models
LLMs are pre-trained on substantial amounts of textual data and have become a significant component of contemporary NLP research. Recent advancements in NLP have led to the development of highly capable LLMs, such as GPT-3 [60], ChatGPT, and GPT-4, which exhibit exceptional performance across a diverse array of NLP tasks, including machine translation, text summarization, and question answering. Concurrently, several previous studies have indicated that LLMs can achieve remarkable results in relevant downstream tasks with minimal or even no demonstration in the prompt [61, 62, 63, 64, 25]. Sánchez et al. [65] proposes a novel neural language model that incorporates inductive biases to enforce explicit relational structures. of pretrained language models. This provides further evidence of the robustness and generality of LLMs.
@@ A.2 ChatGPT & GPT-4
ChatGPT, an advanced LLM developed by OpenAI, is primarily designed for engaging in human-like conversations. During the fine-tuning process, ChatGPT utilizes RLHF [33], thereby enhancing its alignment with human preferences and values.
@@
With the popularity of LLMs, an increasing number of researchers are exploring the specific emergent capabilities and advantages they possess [66]. Bang et al. [62] performs the in-depth analysis of ChatGPT on the multitask, multilingual and multimodal aspects. The findings indicate that ChatGPT excels at zero-shot learning across various tasks, even outperforming fine-tuned models in certain cases. However, it faces challenges when generalized to low-resource languages. Furthermore, in terms of multi-modality, compared to more advanced vision-language models, the capabilities of ChatGPT remain fundamental. Moreover, ChatGPT has garnered considerable attention in other various domains, including information extraction [53, 25], reasoning [29], text summarization [67], question answering [68] and machine translation [69], showcasing its versatility and applicability in the broader field of natural language processing.
While there is a growing body of research on ChatGPT, investigations into GPT-4 continue to be relatively limited. Nori et al. [63] conducts an extensive assessment of GPT-4 on medical competency examinations and benchmark datasets and shows that GPT-4, without any specialized prompt crafting, surpasses the passing score by over 20 points. Kasai et al. [70] also studies GPT-4’s performance on the Japanese national medical licensing examinations. Furthermore, there are some studies on GPT-4 that focus on cognitive psychology [71], academic exams [72], and translation of radiology reports [73].
A.3 LLMs for KG
Now many studies leverage large language models to facilitate the construction of knowledge graphs[74]. Some of these tasks focus on specific subtasks within KG construction. For instance, LLMs are utilized for named entity recognition and classification [75, 76], leveraging their contextual understanding and linguistic knowledge. Furthermore, LLMs have also demonstrated utility in tasks such as relation extraction [25, 77] and link prediction [78, 79, 80]. In line with our approach, several studies have explored the use of LLMs as knowledge bases [81, 82, 83, 74] to support KG construction. For example, some researchers [84] propose a symbolic knowledge distillation framework that extracts symbolic knowledge from LLMs. They first extract commonsense facts from large LLMs like GPT-3, fine-tune smaller student LLMs, and then use these student models to generate KGs. Concurrently, ChatRule [85] uses LLMs to mine logical rules from KGs, addressing computational intensity and scalability issues present in existing methods. ChatRule generates rules with LLMs, integrating the semantic and structural information of KGs, and employs a rule ranking module to evaluate rule quality. These studies highlight the extensive potential of LLMs in KG construction, promoting the automation and intelligent development of this field.
@@ Appendix B Datasets
Entity, Relation and Event Extraction. DuIE2.0 [39] is a substantial Chinese relationship extraction dataset with more than 210,000 sentences and 48 predefined relationship categories. SciERC [40] is a collection of scientific abstracts annotated with seven relations. Re-TACRED [41], an upgraded version of the TACRED dataset, includes over 91,000 sentences across 40 relations. MAVEN [42] a general-domain event extraction benchmark containing 4,480 documents and 168 event types.
Link Prediction. FB15K-237 [43] is widely used as a benchmark for assessing the performance of knowledge graph embedding model on link prediction, encompassing 237 relations and 14,541 entities. ATOMIC 2020 [44] serves as a comprehensive commonsense repository with 1.33 million inferential knowledge tuples about entities and events.
Question Answering. FreebaseQA [45] is an open-domain QA dataset built on the Freebase knowledge graph, comprising various sourced question-answer pairs. MetaQA [16], expanded from WikiMovies [86], provides a substantial collection of single-hop and multi-hop question-answer pairs, surpassing 400,000 in total.
@@ Appendix C Data Collection of VINE
Using GPT-4 data up to September 2021 as a basis, we select a portion of participants’ responses from two competitions organized by the New York Times as part of our data sources. These competitions include the ”February Vocabulary Challenge: Invent a Word” [^6] https://www.nytimes.com/2022/01/31/learning/february-vocabulary-challenge-invent-a-word.html held in January 2022 and the ”Student Vocabulary Challenge: Invent a Word” [^7] https://www.nytimes.com/2023/02/01/learning/student-vocabulary-challenge-invent-a-word.html conducted in February 2023. Both competitions aim to promote the creation of distinctive and memorable new words that address gaps in the English language.
@@ Appendix D Prompts for Evaluation
Event Detection The list of event types: [……]
@@
Link Prediction predict the tail entity [MASK] from the given (40th Academy Awards, time event locations, [MASK]) by completing the sentence ”what is the locations of 40th Academy Awards? The answer is”. predict the tail entity [MASK] from the given (1992 NCAA Men’s Division I Basketball Tournament, time event locations, [MASK]) by completing the sentence ”what is the locations of 1992 NCAA Men’s Division I Basketball Tournament? The answer is ”.The answer is Albuquerque, so the [MASK] is Albuquerque.
predict the tail entity [MASK] from the given (40th Academy Awards, time event locations, [MASK]) by completing the sentence ”what is the locations of 40th Academy Awards? The answer is”.
Question Answering Please answer the following question. Note that there may be more than one answer to the question.
@@
Answer: Android — City Limits